OpenStudy (anonymous):

The solutions to a quadratic equation are x = 2 and x = 7. If k is a nonzero constant, which of the following must be equal to the quadratic equation?

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

a. k(x-2)(x-7)=0 b. (x-2)(x-7)+k=0 c. (x+2)(x+7)+k=0 d. k(x+2)(x+7)=0

OpenStudy (anonymous):

a